OpenStudy (gabylovesyou):

Given that f(x) = 3x + 1 and g(x) = (4x + 2) / 3, solve for g(f(0)). 2 6 8 9

OpenStudy (gabylovesyou):

@ganeshie8

OpenStudy (anonymous):

f(0)=1 and now you solve for g(1)

OpenStudy (anonymous):

you get 4+2=6

OpenStudy (gabylovesyou):

My answer is 2.. am i right ?

OpenStudy (anonymous):

no it is 6

OpenStudy (gabylovesyou):

but dont i have to solve it like this: g(3*0 + 1) = (4(3*0 + 1) + 2) / 3

ganeshie8 (ganeshie8):

2 is right

OpenStudy (gabylovesyou):

yaaay me :)

OpenStudy (gabylovesyou):

Thank you !

ganeshie8 (ganeshie8):

good job!

OpenStudy (anonymous):

oh pellet i didnt see /3 i am really sorry yes so it makes it 6/3=2 sorryy